# Structure of Kerr black hole spacetimes in Weyl conformal gravity

### v1

@@ 3.2 Classification of horizons and ergosurfaces
Both cosmological and Cauchy horizons can be thought of as being generated by an effective repulsive effect. Cosmological horizons generate a background $\mathrm{S}^{+}$ region. Meanwhile, Cauchy horizons can be thought to carve out a $\mathrm{T}$ region within an otherwise fully $\mathrm{S}^{-}$ region. This repulsion is how charge and spin in GR Reissner-Nordstrom and GR Kerr black holes generate an inner $\mathrm{T}$ region not present in GR Schwarzschild black holes.
@@
We similarly label the three ergosurfaces by relation to these horizons. We have the ergosurface $\mathcal{E}_{0}$ at $r^{\mathcal{E}}_{0}=0$ mentioned earlier, the outer ergosurface $\mathcal{E}_{\mathrm{O}}$ found outside the outer horizon $\mathrm{H}_{\mathrm{O}}$, and the cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ located before the cosmological horizon $\mathrm{H}_{\mathrm{C}}$.
We then define Naked singularity spacetimes as spacetimes that do not possess an event horizon. In such naked singularity spacetimes, the only type of horizon that may be present, if any, is a cosmological one.
@@
4 Extremal limits of the CG Kerr spacetimes
@@ 4.1 Extremal horizon limits
Our parameter space of $\gamma,\kappa,$ and $a$ permits three different extremal limits. The first of these is found when $\gamma=\frac{2}{3}$, for any values of $\kappa$ and $a$. This is due to the vanishing of the coefficient $(2-3\gamma)$, giving rise to four repeated roots of the quartic $\Delta^{\mathrm{H}}$ (9).
This particular extremal limit then has four horizons at $r=0$, the only time we have a fourth horizon for $r\geq 0$. We shall refer to this as the Empty case, as we have no horizons for all $r>0$. This is a rather peculiar spacetime, as $\Delta^{\mathrm{H}}=0$ at $r=0$, making the singularity null. We then have $\Delta^{\mathrm{H}}<0$ for all $r>0$, making all of it spacelike $(\mathrm{S})$. As this spacetime does not have a $\mathrm{T}$ region within $r\geq 0$, we are not able to use the sign of $-V^{\prime}_{+}(r)$ (14) to distinguish whether this is $\mathrm{S}^{+}$ or $\mathrm{S}^{-}$. However, as we shall see later on, the Empty case is sandwiched between spacetimes in the parameter maps that only have cosmological horizons, so we can identify this as $\mathrm{S}^{+}$.
We are inclined to consider this particular spacetime as not being well-defined, and may be regarded as a degeneracy in the model, as $\gamma=\frac{2}{3}$ clearly leads to an infinity in $k$ (7). It may be that a sufficient coordinate transformation or conformal transformation can extricate this infinity, but we have been unable to do so here.
The second case is the extremal spin limit $\mathrm{H_{\mathrm{ex(s)}}}$, where the inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$ and outer event horizon $\mathrm{H}_{\mathrm{O}}$ of the black hole merge $(r^{\mathrm{H}}_{\mathrm{I}}=r^{\mathrm{H}}_{\mathrm{O}}\equiv r^{\mathrm{H}}_{\mathrm{ex(s)}})$. For the GR Kerr spacetime, this limit occurs when the mass equals the spin $(a=1)$. For GR Kerr-dS and GR Kerr-AdS spacetimes, the extremal spin value would take on other values $(a\neq 1)$ for different values of $\Lambda$ [56].
Thirdly, we have the extremal horizon cosmological limit $\mathrm{H_{\mathrm{ex(c)}}}$. In this limit, the outer black hole event horizon $\mathrm{H}_{\mathrm{O}}$ and cosmological horizon $\mathrm{H}_{\mathrm{C}}$ coalesce $(r^{\mathrm{H}}_{\mathrm{O}}=r^{\mathrm{H}}_{\mathrm{C}}\equiv r^{\mathrm{H}}_{\mathrm{ex(c)}})$. For GR Kerr-dS, this is known as the rotating Nariai limit [57], as the near-horizon geometry can be shown to reduce to a rotating version of the cosmological Nariai metric [58, 59]. In this work, we do not show that the near-horizon geometry of this limit in CG Kerr spacetimes reduces to a CG corollary of the rotating Nariai spacetime. Therefore, we shall not use the term rotating Nariai here.
@@
Now, when we set the values of two parameters, (21) allows us to find the corresponding values of the third parameter that gives the extremal limits. For example, by setting specific values of $a$ and $\kappa$, we may solve for the extremal values $\gamma^{\mathrm{H}}_{\text{ex}}$. This may give us multiple values. However, when we apply the restriction in (20), we shall get a maximum of two. The larger of these two will give the extremal spin case, and the smaller one the extremal horizon cosmological case $(\gamma^{\mathrm{H}}_{\text{ex(c)}}<\gamma^{\mathrm{H}}_{\text{ex(s)}})$. This reflects the fact that when we plot the locations of horizons in parameter maps of $\gamma$ for set values of $(\kappa,a)$, as we shall do in Section 5, the extremal spin limit is found at a maximum, and the extremal horizon cosmological limit at a minimum.
@@ 4.2 Extremal ergosurface limits
As the ergosurface quartic $\Delta^{\mathcal{E}}$ (8) shows no spin $a$ dependence, we do not have a corollary extremal spin limit for ergosurfaces on the equatorial plane. This is also reflected in the fact that the innermost ergosurface $\mathcal{E}_{0}$ is always at $r^{\mathcal{E}}_{0}=0$.
@@
A cubic with real coefficients may have one or three real roots. The extremal ergosurface cosmological limit corresponds to where two of three real roots are repeated. The condition for this [61] is that the discriminant of $\Delta^{\mathcal{E}}$ (8)
@@
which precisely corresponds to equation (12) in [22], giving the extremal horizon cosmological limit for the CG Schwarzschild case. This correspondence between the extremal horizon cosmological limit in the CG Schwarzschild case and the extremal ergosurface cosmological limit in the rotating CG Kerr case makes sense as $\Delta^{\mathcal{E}}$ (8) has no dependence on the spin $a$. Of course, in the non-rotating CG Schwarzschild spacetimes, the horizons and ergosurfaces are identical.
@@
While we do not have a closed form expression for the $\gamma$ value at the extremal horizon cosmological limit $\mathrm{H}_{\mathrm{ex(c)}}$, the similarity in the forms of $\Delta^{\mathcal{E}}$ (8) and $\Delta^{\mathrm{H}}$ (9) indicate that (26) may serve as a good approximation of it for small enough spin $a$.
@@ 5 General structure and role of $\gamma$
Table 1: Domains of the parametric maps of spacetime features and regions that may be found for $\kappa\geq 0$. For $\kappa<0$, we see these spacetimes supplemented additional domains described in table 2. Here, $\gamma^{\mathrm{H}}_{\mathrm{ex(s)}}$, $\gamma^{\mathrm{H}}_{\mathrm{ex(c)}}$, and $\gamma^{\mathcal{E}}_{\mathrm{ex(s)}}$ are the values of $\gamma$ at the extremal horizon spin limit, extremal horizon cosmological limit, and extremal ergosurface cosmological limit respectively.
@@ I $\gamma>\gamma_{\text{max}}$ $\mathcal{E}_{0}$ $\mathrm{T}$ $\mathrm{N}$
Now, we briefly describe the key features of each domain. Black hole spacetimes are described by the domains $\mathrm{H}_{\mathrm{ex(s)}}$, IV, $\mathcal{E}_{\mathrm{ex(c)}}$, V, and $\mathrm{H}_{\mathrm{ex(c)}}$. All other domains are naked singularity spacetimes, or the degenerate Empty case. We find that the singularity is timelike in all domains, except the Empty case where it is null.
Starting with the horizonless Domain I, where $\gamma>\gamma_{\text{max}}$. We see that the only feature present is the ergosurface $\mathcal{E}_{0}$ at $r=0$. The whole of $r>0$ is both timelike $(\mathrm{T})$ and a non-ergoregion $(\mathrm{N})$. We thus have naked singularities.
@@
We thus find $\gamma_{\text{max}}=1$ from $k=k_{0}=0$. The value of $\gamma_{\text{max}}$ would of course vary when we include $\kappa$.
Thus, as Domain I lies above $\gamma_{\text{max}}$ such that $k<0$, we may consider Domain I as having an Anti-de Sitter background.
In Domain II then, we see the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ and ergosurface $\mathcal{E}_{\mathrm{C}}$ curves asymptotically approach $\gamma_{\text{max}}=1$. Here, $\mathcal{E}_{\mathrm{C}}$ is the boundary for the transition in ergoregion structure $\mathrm{N}\rightarrow\mathrm{E}$ as $r$ is increased. Likewise, the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ demarcates the transition from a timelike region to a cosmological spacelike region $(\mathrm{T}\rightarrow\mathrm{S}^{+})$. At precisely $\gamma=\gamma_{\text{max}}$, these cosmological features are found at $r=+\infty$. As the black hole event horizon is absent, the singularity in these spacetimes is naked.
@@
We find the bulk of the black hole spacetimes within Domain IV, for $\gamma$ values between the extremal spin $\gamma^{\mathrm{H}}_{ex(s)}$ and extremal ergosurface cosmological $\gamma^{\mathcal{E}}_{ex(c)}$ limits. Domain IV shares the causal and ergoregion structure of the extremal spin case.
The Domain IV spacetimes all have a Kerr black hole with a $\mathrm{T}$ region encased by the inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$, followed by an $\mathrm{S}^{-}$ region between $\mathrm{H}_{\mathrm{I}}$ and the outer event horizon $\mathrm{H}_{\mathrm{O}}$. Past the outer horizon, we have another timelike region. Meanwhile, the ergoregion structure is $\mathrm{E}$ between $\mathcal{E}_{0}$ and the outer ergosurface $\mathcal{E}_{\mathrm{O}}$ $(r^{\mathcal{E}}_{0}<r<r^{\mathcal{E}}_{\mathrm{O}})$, followed by $\mathrm{N}$ outside $\mathcal{E}_{\mathrm{O}}$.The cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ and horizon $\mathrm{H}_{\mathrm{C}}$ are then found further out. These demarcate transitions $\mathrm{N}\rightarrow\mathrm{E}$ and $\mathrm{T}\rightarrow\mathrm{S}^{+}$ respectively.
@@ 5.1 Role of $\gamma$
The role of $\gamma$ is much simpler to ascertain in the CG Schwarzschild case (3) when considering the relation between the lapse function $B(r)$ (5) and the potential $\Phi(r)$ is $B(r)=1+2\Phi(r)$. We then have an effective force $\mathbf{f}=-\mathbf{\nabla\phi}$, so for large $r$, $\gamma$ is attractive when $\gamma>0$ and repulsive when $\gamma<0$. This is of course from the $\gamma r$ term in the potential. Thus, in the $\kappa=0$ case of the CG Schwarzschild solution, cosmological horizons are found for $\gamma<0$, while none are found for $\gamma>0$ [22].
We also note that compared to the GR Schwarzschild case, the $-1/r$ term, dominating small $r$ behavior, in the CG Schwarzschild potential is modified by a factor of $(2-3\gamma)$ to become $-(2-3\gamma)/r$. This then switches sign to become positive when $\gamma>\frac{2}{3}$. However, as this becomes relevant at low $r$, the repulsive effect of this generates an inner Cauchy horizon instead of a cosmological horizon. Thus, even as the attractive nature of $\gamma>0$ dominates the large $r$ behavior, when $\kappa=0$, an inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$ may still be present behind an outer event horizon $\mathrm{H}_{\mathrm{O}}$. This is akin to the presence of the inner Cauchy horizon in GR Reissner-Nordstrom-AdS and GR Kerr-AdS, from the repulsive effect of charge and spin respectively, despite an attractive curvature $(\Lambda<0)$ dominating the large $r$ behavior.
Dependence on $\gamma$ is less straightforward when we seek to solve $\Delta^{\mathrm{H}}=0$ (9) in the CG Kerr case. As we can see, no terms are dependent on merely the sign of $\gamma$ itself, but on more complicated combinations of it. For instance, the linear term in this quartic is proportional to $-(2-3\gamma)$. The effect of this term then is dependent on the overall sign of $(2-3\gamma)$ rather than just $\gamma$.
@@
To clarify this, we consider the case of $\kappa=0$, where $k=k_{0}$ (27). As all other factors are squared, the sign of $k_{0}$ is entirely determined by the sign of $(1-\gamma)$. Since this is always positive for $\gamma<\gamma_{max}=1$, $k_{0}>0$ provides a repulsive effect akin to $\Lambda>0$ in GR Kerr-dS. This thus allows cosmological horizons and regions to be found from Domains II downwards in Figure 4. Therefore, unlike the CG Schwarzschild case [22], cosmological horizons may exist for $\gamma>0$ when $\kappa=0$. In fact, due to the $(2-3\gamma)^{2}$ in the denominator, $k_{0}$ becomes very large as we approach $\gamma=\frac{2}{3}$, explaining the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ curves emanating from $(r,\gamma)=(0,\frac{2}{3})$ in Domains II and III.

### v2

@@ 3.2 Classification of horizons and ergosurfaces
Both cosmological and Cauchy horizons can be thought of as being generated by an effective repulsive effect. Cosmological horizons generate an $\mathrm{S}^{+}$ region exterior (at greater $r)$ to them. Meanwhile, Cauchy horizons can be thought to carve out a $\mathrm{T}$ region within an otherwise fully $\mathrm{S}^{-}$ region. This repulsion is how charge and spin in GR Reissner-Nordstrom and GR Kerr black holes generate an inner $\mathrm{T}$ region not present in GR Schwarzschild black holes.
@@
We similarly label the three ergosurfaces by relation to these horizons. We have the ergosurface $\mathcal{E}_{0}$ at $r^{\mathcal{E}}_{0}=0$ mentioned earlier, the outer ergosurface $\mathcal{E}_{\mathrm{O}}$ found outside the outer horizon $\mathrm{H}_{\mathrm{O}}$, and the cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ located interior (at smaller $r$) to the cosmological horizon $\mathrm{H}_{\mathrm{C}}$.
We then define Naked singularity spacetimes as spacetimes with no event horizons $\mathrm{H}_{\mathrm{E}}$ protecting the singularity at the origin. In such naked singularity spacetimes, the only type of horizon that may be present is a cosmological one. This is similar to the naked singularities found in post-extremal GR spacetimes. These violate the Weak Cosmic Censorship Conjecture (WCCC). However, since the WCCC has not been proven in general even in GR [72], we still explore such naked singularity spacetimes here.
@@
4 Extremal limits and horizon temperatures
@@
Furthermore, we compute the surface gravities and Hawking temperatures of horizons in the CG Kerr spacetimes. We then confirm that these quantities vanish at the extremal horizon limits, as would be expected in GR.
@@ 4.1 Extremal horizon limits
Our parameter space of $\gamma,\kappa,$ and $a$ permits three different extremal limits. The first of these is found when $\gamma=\frac{2}{3}$, for any values of $\kappa$ and $a$. This is due to the vanishing of the coefficient $(2-3\gamma)$, giving rise to four coincident roots of the quartic $\Delta^{\mathrm{H}}$ (15).
This particular extremal limit then has four horizons at $r=0$, the only time we have a fourth horizon for $r\geq 0$. We shall refer to this as the Empty case, as we have no horizons for all $r>0$. This is a rather peculiar spacetime, as $\Delta^{\mathrm{H}}=0$ at $r=0$, making the singularity null. We then have $\Delta^{\mathrm{H}}<0$ for all $r>0$, making all of it spacelike $(\mathrm{S})$. As this spacetime has no $\mathrm{T}$ region within $r\geq 0$, we cannot use the sign of $V^{\prime}_{+}(r)$ (20) to distinguish whether this is $\mathrm{S}^{+}$ or $\mathrm{S}^{-}$. However, as we shall see later on, the Empty case is sandwiched between spacetimes in the parameter maps that only have cosmological horizons, so we can identify this as $\mathrm{S}^{+}$.
We are inclined to consider this particular spacetime to be ill-defined. It may be regarded as a degeneracy in the model, as $\gamma=\frac{2}{3}$ clearly leads to an infinity in $k$ (12). If some coordinate or conformal transformation can extricate this infinity, we have failed to discover it.
[…]
@@
[…]
@@ 4.2 Extremal ergosurface limits
As the ergosurface quartic $\Delta^{\mathcal{E}}$ (13) shows no spin $a$ dependence, we have no analogue of the extremal spin limit for ergosurfaces on the equatorial plane. This is also reflected in the fact that the innermost ergosurface $\mathcal{E}_{0}$ is always at $r^{\mathcal{E}}_{0}=0$.
@@
A cubic with real coefficients may have one or three real roots. The extremal ergosurface cosmological limit $\mathcal{E}_{\mathrm{C}}$ corresponds to where two of three real roots coincide. The condition for this [78] is that the discriminant of $\Delta^{\mathcal{E}}$ (13)
@@
which precisely corresponds to equation (12) in [27], giving the extremal horizon cosmological limit for the CG Schwarzschild case. This correspondence between the extremal horizon cosmological limit in the CG Schwarzschild case and the extremal ergosurface cosmological limit in the rotating CG Kerr case makes sense. Since, as discussed earlier in subsection 2.1, we are operating on the equatorial plane $(\theta=\frac{\uppi}{2})$, $\Delta^{\mathcal{E}}$ (13) has no dependence on $a$. Of course, in the non-rotating CG Schwarzschild spacetimes, the horizons and ergosurfaces are identical.
@@
While we have not discovered a closed form expression for the $\gamma$ value at the extremal horizon cosmological limit $\mathrm{H}_{\mathrm{ex(c)}}$, the similarity in the forms of $\Delta^{\mathcal{E}}$ (13) and $\Delta^{\mathrm{H}}$ (15) indicate that (32) may serve as a good approximation of it for small enough spin $a$.
4.3 Horizon temperatures
We now compute the surface gravities and semiclassical Hawking temperatures [56, 57] of the horizons found in the CG Kerr spacetimes.
As outlined in [79], for a metric in Boyer-Lindquist coordinates, of the form (11), the surface gravity of the black hole event horizon $\mathrm{H}_{\mathrm{O}}$ is given by
$K_{\mathrm{O}}=\frac{1}{2}\frac{\partial_{r}\left[\Delta^{\mathrm{H}}\right]}{r^{2}+a^{2}}\ \Bigg{|}_{r=r^{\mathrm{H}}_{\mathrm{O}}}=\frac{1}{2}\frac{-4k(r^{\mathrm{H}}_{\mathrm{O}})^{3}+2(r^{\mathrm{H}}_{\mathrm{O}})-2\widetilde{M}}{(r^{\mathrm{H}}_{\mathrm{O}})^{2}+a^{2}},$ (33)
where $k$ and $\widetilde{M}$ are as defined in (12). From this, the Hawking temperature $T$ of this event horizon is then
$T_{\mathrm{O}}=\frac{K_{\mathrm{O}}}{2\uppi}=\frac{1}{4\uppi}\frac{-4k(r^{\mathrm{H}}_{\mathrm{O}})^{3}+2(r^{\mathrm{H}}_{\mathrm{O}})-2\widetilde{M}}{(r^{\mathrm{H}}_{\mathrm{O}})^{2}+a^{2}}.$ (34)
Meanwhile, the respective formulas for the repulsive Cauchy $\mathrm{H}_{\mathrm{I}}$ and cosmological $\mathrm{H}_{\mathrm{C}}$ horizons take on an overall negative sign to give
$K_{\mathrm{I/C}}=-\frac{1}{2}\frac{\partial_{r}\left[\Delta^{\mathrm{H}}\right]}{r^{2}+a^{2}}\ \Bigg{|}_{r=r^{\mathrm{H}}_{\mathrm{{I/C}}}}=-\frac{1}{2}\frac{-4k(r^{\mathrm{H}}_{\mathrm{{I/C}}})^{3}+2(r^{\mathrm{H}}_{\mathrm{{I/C}}})-2\widetilde{M}}{(r^{\mathrm{H}}_{\mathrm{{I/C}}})^{2}+a^{2}},$ (35)
and
$T_{\mathrm{I/C}}=\frac{K_{\mathrm{I/C}}}{2\uppi}=-\frac{1}{4\uppi}\frac{-4k(r^{\mathrm{H}}_{\mathrm{{I/C}}})^{3}+2(r^{\mathrm{H}}_{\mathrm{{I/C}}})-2\widetilde{M}}{(r^{\mathrm{H}}_{\mathrm{{I/C}}})^{2}+a^{2}}.$ (36)
[…]
$K_{\mathrm{ex(s)}}=K_{\mathrm{I}}(r^{\mathrm{H}}_{\mathrm{I}}=r^{\mathrm{H}}_{\mathrm{ex(s)}})\ +K_{\mathrm{O}}(r^{\mathrm{H}}_{\mathrm{O}}=r^{\mathrm{H}}_{\mathrm{ex(s)}})=0\rightarrow T_{\mathrm{ex(s)}}=0$ (37)
$K_{\mathrm{ex(c)}}=K_{\mathrm{O}}(r^{\mathrm{H}}_{\mathrm{O}}=r^{\mathrm{H}}_{\mathrm{ex(c)}})+K_{\mathrm{C}}(r^{\mathrm{H}}_{\mathrm{C}}=r^{\mathrm{H}}_{\mathrm{ex(c)}})=0\rightarrow T_{\mathrm{ex(c)}}=0.$ (38)
We have thus shown that the horizon surface gravities $K$ and Hawking temperatures $T$ both vanish at the extremal horizon limits in CG Kerr spacetimes, just as we would expect for the GR Kerr and GR Kerr-(A)dS extremal horizon limits.
@@ 5 General structure and role of $\gamma$
Table 1: Domains of the parametric maps of spacetime features and regions that may be found for $\kappa\geq 0$. For $\kappa<0$, we see these spacetimes supplemented by additional domains described in table 2. Here, $\gamma^{\mathrm{H}}_{\mathrm{ex(s)}}$, $\gamma^{\mathrm{H}}_{\mathrm{ex(c)}}$, and $\gamma^{\mathcal{E}}_{\mathrm{ex(s)}}$ are the values of $\gamma$ at the extremal horizon spin limit, extremal horizon cosmological limit, and extremal ergosurface cosmological limit respectively.
@@ I $\gamma>\gamma_{\text{max}}$ $\mathcal{E}_{0}$ $\mathrm{T}$ $\mathrm{N}$
Now, we briefly describe the key features of each domain. Black hole spacetimes, containing an event horizon $\mathrm{H}_{\mathrm{C}}$, are described by the domains $\mathrm{H}_{\mathrm{ex(s)}}$, IV, $\mathcal{E}_{\mathrm{ex(c)}}$, V, and $\mathrm{H}_{\mathrm{ex(c)}}$. All other domains are naked singularity spacetimes, or the degenerate Empty case. We find that the singularity is timelike in all domains, except the Empty case where it is null.
Starting with the horizonless Domain I, where $\gamma>\gamma_{\text{max}}$. We see that the only feature present is the ergosurface $\mathcal{E}_{0}$ at $r=0$. The whole of $r>0$ is both timelike $(\mathrm{T})$ and a non-ergoregion $(\mathrm{N})$. Domain I thus has a naked singularity at $r=0$.
@@
We thus find $\gamma_{\text{max}}=1$ from $k=k_{0}=0$. The value of $\gamma_{\text{max}}$ would of course vary when we include $\kappa$. Thus, as Domain I lies above $\gamma_{\text{max}}$ such that $k<0$, we may consider Domain I as having an Anti-de Sitter background.
In Domain II then, we see the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ and ergosurface $\mathcal{E}_{\mathrm{C}}$ curves asymptotically approach $\gamma_{\text{max}}=1$. Here, $\mathcal{E}_{\mathrm{C}}$ is the boundary for the transition in ergoregion structure $\mathrm{N}\rightarrow\mathrm{E}$ as $r$ is increased. Likewise, the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ demarcates the transition from a timelike region to a cosmological spacelike region $(\mathrm{T}\rightarrow\mathrm{S}^{+})$. At precisely $\gamma=\gamma_{\text{max}}$, these cosmological features ($\mathrm{H}_{\mathrm{C}}$ and $\mathcal{E}_{\mathrm{C}}$) are found at $r=+\infty$. As the black hole event horizon is absent, the singularity in these spacetimes is naked.
@@
We find the bulk of the black hole spacetimes within Domain IV, for $\gamma$ values between the extremal spin $\gamma^{\mathrm{H}}_{ex(s)}$ and extremal ergosurface cosmological $\gamma^{\mathcal{E}}_{ex(c)}$ limits. Domain IV shares the causal and ergoregion structure of the extremal spin case. The Domain IV spacetimes all have a Kerr black hole with a $\mathrm{T}$ region encased by the inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$, followed by an $\mathrm{S}^{-}$ region between $\mathrm{H}_{\mathrm{I}}$ and the outer event horizon $\mathrm{H}_{\mathrm{O}}$. Past the outer horizon, we have another timelike region. Meanwhile, the ergoregion structure is $\mathrm{E}$ between $\mathcal{E}_{0}$ and the outer ergosurface $\mathcal{E}_{\mathrm{O}}$ $(r^{\mathcal{E}}_{0}<r<r^{\mathcal{E}}_{\mathrm{O}})$, followed by $\mathrm{N}$ outside $\mathcal{E}_{\mathrm{O}}$.The cosmological ergosurface $\mathcal{E}_{\mathrm{C}}$ and horizon $\mathrm{H}_{\mathrm{C}}$ are then found further out. These demarcate transitions $\mathrm{N}\rightarrow\mathrm{E}$ and $\mathrm{T}\rightarrow\mathrm{S}^{+}$ respectively.
@@ 5.1 Role of $\gamma$
We see that the dependence of the CG Kerr solution  (11) on $\gamma$, as seen in figure 4, is rather complicated. We thus first turn to the CG Schwarzschild case (8), where this dependence on $\gamma$ is more straightforward by considering that the relation between the lapse function $B(r)$ (10) and the potential $\Phi(r)$ is $B(r)=1+2\Phi(r)$. We then have an effective force $\mathbf{f}=-\mathbf{\nabla\Phi}$, so that for large $r$, the linear $\gamma r$ term in the potential is attractive when $\gamma>0$ and repulsive when $\gamma<0$. Thus, in the $\kappa=0$ case of the CG Schwarzschild solution, cosmological horizons are found for $\gamma<0$, while none are found for $\gamma>0$ [27].
We also note that compared to the GR Schwarzschild case, the 1/r-1/r term, dominating small $r$ behavior, in the CG Schwarzschild potential is modified by a factor of $(2-3\gamma)$ to become $(2-3\gamma)/r$. This then switches sign to become positive when $\gamma>\frac{2}{3}$. However, as this becomes relevant at low $r$, the repulsive effect of this generates an inner Cauchy horizon instead of a cosmological horizon. Thus, even as the attractive nature of the $\gamma r$ term for $\gamma>0$ dominates the large $r$ behavior, when $\kappa=0$, an inner Cauchy horizon $\mathrm{H}_{\mathrm{I}}$ may still be present behind an outer event horizon $\mathrm{H}_{\mathrm{O}}$. This is akin to the presence of the inner Cauchy horizon in GR Reissner-Nordstrom-AdS and GR Kerr-AdS, from the repulsive effect of charge and spin respectively, despite an attractive curvature $(\Lambda<0)$ dominating the large $r$ behavior.
Dependence on $\gamma$ is less straightforward when we seek to solve $\Delta^{\mathrm{H}}=0$ (15) in the CG Kerr case. As we can see in  (15), no terms depend on merely the sign of $\gamma$ itself, but on more complicated combinations of it. For instance, the linear term in this quartic is proportional to $(2-3\gamma)$. The effect of this term then depends on the sign of $(2-3\gamma)$ rather than just $\gamma$.
@@
To clarify this, we consider the case of $\kappa=0$, where $k=k_{0}$ (39). As all other factors are squared, the sign of $k_{0}$ is entirely determined by the sign of $(1-\gamma)$. Since this is always positive for $\gamma<\gamma_{max}=1$, $k_{0}>0$ provides a repulsive effect akin to $\Lambda>0$ in GR Kerr-dS. This thus allows cosmological horizons and $S^{+}$ regions to be found from Domains II downwards in figure 4. Therefore, unlike the CG Schwarzschild case [27], cosmological horizons may exist for $\gamma>0$ when $\kappa=0$. In fact, due to the $(2-3\gamma)^{2}$ in the denominator, $k_{0}$ becomes very large as we approach $\gamma=\frac{2}{3}$, explaining the cosmological horizon $\mathrm{H}_{\mathrm{C}}$ curves emanating from $(r,\gamma)=(0,\frac{2}{3})$ in Domains II and III.